OpenStudy (anonymous):

A boat sails along the shore. To an observer, the boat appears to move at a speed of 3 m/s, and a man on the boat walking forward appears to have a speed of 4.2 m/s. How long does it take the man to move 8 m relative to the boat?

OpenStudy (anonymous):

If I read this right,the man is walking forward at 4.2m/s relative to an observer on shore and the boat is moving at 3 m/s relative to the shore.. therefore the man is walking forward at 1.2 m/s relative to the BOAT . Using the displacement of 8 m forward and the relative velocity of 1.2m/s forward, you can then use t=d/v to get your answer. ( 8m/1.2 m/s= your answer).

OpenStudy (anonymous):

4.2-3=1.2*8=9.6seconds

OpenStudy (anonymous):

6.7 seconds